x^(-3/5)/x^(-4/3)=150
Answer by lwsshak3(11628)

x^(-3/5)/x^(-4/3)=150
x^[(-3/5)-(-4/3)]=150
x^[(-9/15)-(-20/15)]=150
x^[(-9/15)-(-20/15)]=150
x^(11/15)=150
raise both sides to 15/11
x=150^(15/11)≈927.6778..